Find the total amount of money received if Tshs 800000 is deposited in a bank at the rate of 9% compounded semi annually in one year
Answer by Theo(13342)

the interest rate is 9% per year compounded semi-annually.
divide 9% by 2 to get 4.5% every semi-annual period.
800,000 * 1.045 ^ 2 = 873,620.
her nominal interest rate per year was 9%.
her effective interest rate per year was 1.045 ^ 2 = 1.092025 - 1 = .092025 * 100 = 9.2025%.
873620 / 800000 = 1.092025 - 1 = .092025 * 100 = 9.2025%.
your solution is 873,620.
that's how much she will have at the end of the year after investing 800000 at 9% compounded semi-annually for that one year.